OpenStudy (tylerbanks2001):

Jason estimates that his car loses 12% of its value every year. The initial value is $12,000. Which best describes the graph of the function that represents the value of the car after x years? A.f(x) = 12,000(0.12)x, with a horizontal asymptote of y = 0 B.f(x) = 12,000(1.12)x, with a vertical asymptote of x = 0 C.f(x) = 12,000(0.88)x, with a horizontal asymptote of y = 0 D.f(x) = (12,000 0.88)x, with a vertical asymptote of x = 0

OpenStudy (anitas.):

wut do you think it is?

OpenStudy (anitas.):

@TylerBanks2001

OpenStudy (tylerbanks2001):

um idk thats why i asked

OpenStudy (anitas.):

the form of the equation is y=starting-value* ( 1 - loss%) ^ x y=12000* ( 1 - 12%) ^ x y=12000* ( 1 - .12) ^ x y=12000* ( .88) ^ x so which one is it from your list?